OpenStudy (anonymous):

simplify the equation. show your work. (x – 2) (3x-4)

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

isn't it already factored?

OpenStudy (anonymous):

not sure/ these problems confuse me

OpenStudy (anonymous):

@Bill_Cipher

OpenStudy (anonymous):

@alexandrea.lane

OpenStudy (anonymous):

yeah i'm not really sure how much farther you can factor this.

OpenStudy (anonymous):

then can you help me with another one

OpenStudy (anonymous):

A water balloon is tossed into the air with an upward velocity of 25 ft/s. Its height h(t) in ft after t seconds is given by the function h(t) = − 16t^2 + 25t + 3. Show your work.

OpenStudy (anonymous):

after how many seconds will the balloon hit the ground? use quadratic formula

OpenStudy (anonymous):

I'm not sure if this would help but (x-2)(3x-4)=1/3(3x-5)^2-1/3=x(3x-10)+8

OpenStudy (anonymous):

that was for the previous question

OpenStudy (anonymous):

(-b+sqrt(b^2-4ac)/2a

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

a=-16, b=25, c=3 That was the quadratic formula above

OpenStudy (anonymous):

substitute the values for a, b, and c into the formula

OpenStudy (anonymous):

-25+sqrt+(25^2-4*-16*3)/2*-16

OpenStudy (anonymous):

yep now simplify

OpenStudy (anonymous):

how

OpenStudy (anonymous):

25^2-4*16*3=?

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

817

OpenStudy (anonymous):

so 817 seconds

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

finish the rest of the formula right now you have (-25+sqrt(817))/-32

OpenStudy (anonymous):

or actually in this case it would actually be (-25-sqrt(817))/-32

OpenStudy (anonymous):

can you help

OpenStudy (anonymous):

or (25+sqrt(817))/32

OpenStudy (anonymous):

or you could graph the equation here at https://www.desmos.com/ to find the answer

OpenStudy (anonymous):

i dont need a graph though

OpenStudy (anonymous):

can you just help me?

OpenStudy (anonymous):

did you look at what I side before about the quadratic formula?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay. Tell me what you were confused on and I will help you

OpenStudy (anonymous):

sqrt what is that and how do i solve

OpenStudy (anonymous):

square root

OpenStudy (anonymous):

so can you help me finish it?

OpenStudy (anonymous):

okay so what is the quadratic formula

OpenStudy (anonymous):

(-b+sqrt(b^2-4ac)/2a

OpenStudy (anonymous):

okay and what are a, b, and c

OpenStudy (anonymous):

a=-16, b=25, c=3

OpenStudy (anonymous):

so plug in these values (into the quadratic formula)

OpenStudy (anonymous):

i did and you aid i had to finish it out and so far i have -25+sqrt+817/-32

OpenStudy (anonymous):

said

OpenStudy (anonymous):

so how many seconds

OpenStudy (anonymous):

in this case since the result needs to be positive so the answer would be (25+sqrt(817))/32

OpenStudy (anonymous):

that's how many seconds

OpenStudy (anonymous):

that doesnt make sense

OpenStudy (anonymous):

sorry if I was so confusing but it said to use quadratic formula so it probably wasn't going to be a whole number

OpenStudy (anonymous):

(25+sqrt(817))/32

OpenStudy (anonymous):

thats the amount of seconds?

OpenStudy (anonymous):

yep if you graph it on desmos.com you will see that this is wrong

OpenStudy (anonymous):

i mean right

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

What will the height be at t = 1 second

OpenStudy (anonymous):

this is the question that follows this one on the worksheet

OpenStudy (anonymous):

can you help with my final question?

OpenStudy (anonymous):

yeah okay

OpenStudy (anonymous):

just plug in 1 into the equation -16t^2+25t+3

OpenStudy (anonymous):

-16*1^2+25*1+3

OpenStudy (anonymous):

now simplify

OpenStudy (anonymous):

-16*25+3

OpenStudy (anonymous):

-397

OpenStudy (anonymous):

you mean -16+25+3=12

OpenStudy (anonymous):

how

OpenStudy (anonymous):

look at your above post (above when I said "now simplify")